OpenStudy (anonymous):

y=4x+3 = ? ordered pairs ,please help

Directrix (directrix):

Welcome to Open Study.

Directrix (directrix):

y=4x+3 is the equation of a line. A line is made up of infinitely many points. To find coodinates of points on a given line, select a value of x of your choosing. Then, compute the y value that goes with the value of x you chose.

Directrix (directrix):

y=4x+3 ---------- Example. Let x = -2. Then, y = ? y = 4x + 3 y = 4 (-2) + 3 y = -8 + 3 y = -5 The ordered pair is (-2,-5)